Hulk is helping Natasha pull a 67 - kg box of weapons up a39° inclined plane with a velocity of 6.5 m/s. Hulk is pulling it with a force of 574 N parallel to the 6.19 m incline. The coefficient of kinetic friction is 0.39. What is the speed of the crate at the end of the incline?


1
Expert's answer
2021-03-03T11:26:08-0500

Let's first apply the Newton's Second Law of Motion in projections on axis "x" and "y":


"F_{pull}-mgsin\theta-F_{fr}=ma,""N-mgcos\theta=0."

Let's rewrite our equations:


"F_{pull}-mgsin\theta-\mu_kN=ma,""N=mgcos\theta."

Let's substitute "N" into the previous equation and find the acceleration of the box:


"a=\dfrac{F_{pull}-mgsin\theta-\mu_kmgcos\theta}{m}=\dfrac{F_{pull}-mg(sin\theta+\mu_kcos\theta)}{m},""a=\dfrac{574\ N-67\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot(sin39^{\circ}+0.39\cdot cos39^{\circ})}{67\ kg},""a=0.804\ \dfrac{m}{s^2}."

Finally, we can find the final speed of the box at the end of the incline from the kinematic equation:


"v_f^2=v_0^2+2as,""v_f=\sqrt{v_0^2+2as},""v_f=\sqrt{(6.5\ \dfrac{m}{s})^2+2\cdot0.804\ \dfrac{m}{s^2}\cdot6.19\ m}=7.22\ \dfrac{m}{s}."